OpenStudy (anonymous):

I need URGENT help with factoring 2a^2+30a+100....Please help me I have tried everything:(

OpenStudy (anonymous):

2a^2+30a+100=2a^2+20a+10a+100=2a(a+10)+10(a+10)=(2a+10)(a+10)

OpenStudy (anonymous):

Okay thanks but like how did you do that??

OpenStudy (anonymous):

first multiply 2 and 100 then find two numbers whose product is 2 X 100=200 and sum is 30

OpenStudy (anonymous):

Ohhh I see ugh Im such a blonde:)

OpenStudy (precal):